# Isoparametric functions on Finsler space forms $*{}^{*}\,$

### v2

@@
Abstract: In this paper, we prove that transnormal functions are isoparametric functions on a Finsler space form $(N(c),F)$ under certain conditions, which generalize Theorem B in [5]. Further more, we obtain a local isoparametric function constructed by a distance function and give a global isoparametric function on a standard Finsler sphere.
@@ 1 Introduction
The study on isoparametric theory has a long history. A smooth function $f$ defined on a connected complete Riemannian manifold $(N,g)$ is called an isoparametric function if there exists a smooth function $a$ and a continuous function $b$ satisfying
[…]
E. Cartan began to study isoparametric hypersurfaces in real space forms with constant sectional curvature $c$ systematically [1]. Then many mathematicians started working on isoparametric functions and the classification of isoparametric hypersurfaces in a real space form [2, 3, 4]. If $f$ only satisfies the first equation of (1.1), $f$ is called a transnormal function.
@@
However, no proof was given in [5]. Later,  [6] supplied the proof of Theorem A and constructed an example to point out there is an essential difference between transnormal functions and isoparametric functions in hyperbolic spaces.
@@
[…]
@@
The fundamental structural result in [5] was generalized to Finsler geometry in [13]. Based on the structural result given in [13], in this paper, we try to generalize Theorem A to Finsler geometry. Unlike the Riemannian case, the geodesic are not necessarily reversible and the regularity of $f$, $a(t)$ and $b(t)$ in (1.2) are very poor, which causes more difficulties for the research.
For a connected Finsler manifold $(N,F)$, set $\overline{J}=f(N)$. Denote the focal variety of $f$ by $M_{\pm}$, which means the set where $f$ attains its global maximum value or global minimum value. Let $f$ be a $C^{2}$ transnormal function satisfying $F(\nabla f)^{2}=a(f)$, where $a(t)\neq 0$ for $t\in J$ and $a^{\prime}(t)\neq 0$ for $t\in\partial\overline{J}$. We called such $f$ an appropriate transnormal function. Let $g$ be the number of distinct principal curvatures for an isoparametric hypersurface. Then we have
@@
Let $(N(c),F)$ be a Finsler space form. Suppose that $f$ and $-f$ are both appropriate transnormal functions. Then $f$ is isoparametric when $c\geq 0$, or $c<0$ and all principal curvatures $\kappa_{i}$ of a regular level hypersurface satisfying $|\kappa_{i}|\geq\sqrt{-c}$, $1\leq i\leq g$.
@@
Let $M_{\pm}$ be submanifolds and $f$ has no critical value in $J$. Then $f$ and $-f$ are both appropriate transnormal functions in Theorem 1.1 can be weaken to $f$ and $-f$ are transnormal functions.
@@
Let $(N(c),F)$ be a forward Finsler space form and $r(x)=d_{F}(p,x)$ be a distance function from any $p\in N$. Then $f=\frac{1}{2}r^{2}(x)$ is a local isoparametric function with $g=1$ on the cut-domain of $p$. Particularly, $f=-\cos(\sqrt{c}r)$ is a global isoparametric function on a standard Finsler sphere $(\mathbb{S}^{n}(c),F)$.
@@ 2.1 Finsler manifolds
$\nabla^{w}_{v}X(x)=\{v^{j}\frac{\partial X^{i}}{\partial x^{j}}(x)+\Gamma^{i}_{jk}(w)v^{j}X^{k}(x)\}\frac{\partial}{\partial x^{i}}.$
@@
where $X,Y\in T_{x}N$, $\nabla^{2}f(x)=\nabla^{\nabla f}_{X}\nabla f\in T_{x}^{*}N\otimes T_{x}N$ for $x\in N_{f}$. The Laplacian of $f$ can defined by
@@ 2.2 Anisotropic submanifolds
Let $(N,F)$ be an $n$-dimensional Finsler manifold and $\phi:M\to(N,F)$ be an $(n-1)$-dimensional immersion. Set $(x,y)\in TN$ with local coordinate $(x^{i},y^{i})$ and $(u,v)\in TM$ with local coordinate $(u^{a},v^{a})$, we have
@@
We call the eigenvalues of ${A}_{\textbf{n}}$, $k_{1},k_{2},\cdots,k_{m}$, the principal curvatures of $M$ with respect to n. Define $\hat{H}_{\textbf{n}}=\sum\limits_{a=1}^{m}k_{a}$ the anisotropic mean curvature of $M$.
If $m=n-1$, there exists a global unit normal vector field n. Then $(M,\hat{g})$ is an oriented anisotropic hypersurface and the Weigarten formula (2.8) becomes
@@ 2.3 Isoparametric functions and isoparametric hypersurfaces
More precisely, the function $f$ is called isoparametric (resp. $d\mu$-isoparametric) on $(N,F,d\mu)$, if (1.2) holds for $\Delta f=\hat{\Delta}f$ (resp. $\Delta f=\Delta_{\sigma}f$), which is defined by (2.5) (resp. (2.6)). Each regular level hypersurfaces $M_{t}=f^{-1}(t)$ is called an ($d\mu$)-isoparametric hypersurface on $(N,F,d\mu)$.
If $(N,F,d\mu)$ has constant $\mathbf{S}$-curvature, then $f$ is isoparametric if and only if it is $d\mu$-isoparametric [7]. Therefore, for simplicity, we only consider $\Delta f$ as $\hat{\Delta}f$ in this paper. Obviously, all conclusions in this paper still hold for $d\mu$-isoparametric hypersurfaces if $\mathbf{S}$-curvature is a constant.
If for any $x\in N$, there exists a neighborhood $U$ of $x$ and an isoparametric function $f$ defined on $U$, such that $N\cap U$ is a regular level hypersurface of $f$, then $N$ is called a local isoparametric hypersurface [16].
3 Transnormal functions in Finsler space forms
@@ 3.1 Transnormal functions and focal varieties
[13] Suppose that $\beta$ is the only critical value of a transnormal function $f$ in $[\alpha,\beta]\subset\overline{J}$, then
$d(M_{\alpha},M_{\beta})=\int^{\beta}_{\alpha}\frac{df}{\sqrt{a(f)}}$ (3.1)
and the improper integral in (3.1) converges.
@@
[13] Let $(N,F)$ be a complete, connected Finsler manifold and $f$ be a $C^{2}$ transnormal function on $N$. If $a(t)\neq 0$ for $t\in J$ and $a^{\prime}(t)\neq 0$ for $t\in\partial\overline{J}$, then we have
(1)each focal variety of $f$ is a smooth submanifold of $N$.
From Lemma 3.2 and Lemma 3.5 in [13], we have
Lemma 3.3.
Let $\gamma(s)$ be a maximal $f$-segment satisfying $\gamma(0)\in M_{+}$ (resp. $M_{-}$), then $\gamma^{\prime}(0)\in\mathcal{N}^{0}_{\gamma(0)}M_{+}$ (resp. $\mathcal{N}^{0}_{\gamma(0)}M_{-}$). Conversely, for any $\eta\in\mathcal{N}^{0}_{p}M_{+}$ (resp. $\mathcal{N}^{0}_{q}M_{-}$), there exists a maximal $f$-segment $\gamma(s)$ satisfying $\gamma(0)=p\in M_{+}$ (resp. $\gamma(0)=q\in M_{-}$) and $\gamma^{\prime}(0)=\eta$.
From Lemma 3.3, we know that an $f$-segment and a normal vector on $M_{\pm}$ are one-to-one correspondences. Set $\overline{J}=[\alpha,\beta]$. Since $f$ is an appropriate transnormal function on a complete connected manifold $(N,F)$ , we define $d_{+}=d(M_{-},M_{+})=\int^{\beta}_{\alpha}\frac{dt}{\sqrt{a(t)}}$. Since $-f$ is still an appropriate transnormal function on $(N,F)$, set $F(\nabla(-f))=\tilde{a}(-f)$. Denote $d_{-}=d(M_{+},M_{-})=\int^{\beta}_{\alpha}\frac{dt}{\sqrt{\tilde{a}(t)}}$. In fact, $M_{\pm}^{f}=M_{\mp}^{-f}$. Hence, from Lemma 3.3, we immediately have
Lemma 3.4.
Let $d_{+}<+\infty$ and $\gamma(s)$, $s\in[0,+\infty)$, be a unit speed geodesic such that $\gamma(s)\big{|}_{[0,d_{+}]}$ is a maximal $f$-segment, then $\gamma(s)\big{|}_{[d_{+},d_{+}+d_{-}]}$ is a maximal $(-f)$-segment with respect to $F$.
@@
If $M_{\pm}$ are submanifolds and ±$\pm f$ are transnormal functions with no critical values in $J$, then Lemma 3.3 and Lemma 3.4 hold.
@@
In this case, $d_{+},d_{-}<+\infty$. From Lemma 3.6, for any $k_{i}$, there exists some $s_{i}$ such that $\gamma(s_{i})\in Foc(f)$. Then From Lemma 3.4 and (3.5), there must exist $n_{i},m_{i}\in Z$ such that the principal curvature can be expressed by
@@
From Lemma 3.6, for any $k_{i}\neq 0$, there exists some $s_{i}$ such that $\gamma(s_{i})\in Foc(f)$.
@@
(2) $Foc(f)\neq\varnothing$. From [13], $p=\gamma(r_{0})$ is the focal point defined by the normal exponential map $\Phi(s,t):=E(\sigma(t),s\textbf{n}(\sigma(t)))$, where $\sigma(t)$ is a curve on $M$ and $\textbf{n}(\sigma(t))\in\mathcal{N}^{0}_{\sigma(t)}M$, if and only if $J(r_{0})=\frac{\partial\Phi}{\partial t}\big{|}_{t=0,s=r_{0}}=0$. From Cartan-Hadamard Theorem, $\gamma(s)$ passes through $Foc(f)$ once. Hence, from Lemma 3.4 and (3.5), we have $g\leq 2$. One principal curvature of $M$ is $\frac{1}{r}$ or $\frac{1}{r+d_{-}}$. If $g=2$, the other principal curvature of $M$ is 00. That is, $M$ is an isoparametric hypersurface.
@@
From Cartan-Hadamard Theorem, $Foc(f)=M_{+}$ or $Foc(f)=M_{-}$. Meanwhile, there exists one $k_{i}$ at most such that $|\kappa_{i}|>\sqrt{-c}$. Hence, $g\leq 2$. From Lemma 3.4 and (3.5), the principal curvatures may be $\sqrt{-c}\coth\sqrt{-c}r$, $\sqrt{-c}\coth\sqrt{-c}(r+d_{-})$, $\sqrt{-c}$ or $-\sqrt{-c}$. Hence, $\kappa_{i}$ are constants for any $1\leq i\leq g$.
By case 1, case 2 and case 3, the principal curvatures $\kappa_{i}$ of $M$ are constants. Further more, from (2.10),
@@
which is an ODE with respect to an arc-length parameter $s$. From (3.6), we know each level hypersurface of $f$ has constant principal curvatures. This completes the proof of Theorem 1.1. From the proof of Theorem 1.1, we immediately have
@@
Let $(N(c),F)$ be a Finsler space form. Suppose that $f$ and $-f$ are both appropriate transnormal functions on $N(c)$. If $c=0$, or $c<0$ and all principal curvatures $\kappa_{i}$ of a regular level hypersurface satisfying $|\kappa_{i}|\geq\sqrt{-c}$, $1\leq i\leq g$, then each regular level hypersurface has two distinct principal curvatures at most.
@@
In [12], we obtain the same result in different conditions. In fact, we need to rectify an omission of Theorems in [12]. That is, for $c<0$, it also need to satisfy $|\kappa_{i}|\geq\sqrt{-c}$.
@@
In local case, Theorem 3.7 doesn’t hold. That is, there exist local isoparametric hypersurfaces in some Minkowski spaces with more than two principal curvatures (See Proposition 3.7 for detail).
If $F$ is reversible, $\tilde{a}(-f)=F^{*}(-df)=F^{*}(df)=F(\nabla f)=a(f)$. That is, $f$ is an appropriate transnormal function if and only if $-f$ is an appropriate transnormal function. Hence, we have
@@
4 Local isoparametric hypersurfaces

### v3

@@
Abstract: In this paper, we prove that transnormal functions are isoparametric functions on Finsler space forms $(N(c),F)$ under certain conditions, which generalize Theorem B given by Q. M. Wang in Riemannian case. Next, we discuss the relationship between umbilic hypersurfaces and isoparametric hypersurfaces in $(N(c),F)$. Further more, we construct a locally isoparametric function by a distance function and give a global isoparametric function on a standard Finsler sphere.
@@ 1 Introduction
The study on isoparametric theory has a long history. A smooth function $f$ defined on a connected complete Riemannian manifold $(N,g)$ is called an isoparametric function if there exists a smooth function $\bar{a}(t)$ and a continuous function $\bar{b}(t)$ satisfying
$\left\{\begin{aligned} &\|\nabla f\|^{2}=\bar{a}(f),\\ &\Delta f=\bar{b}(f).\end{aligned}\right$. (1.1)
E. Cartan began to study isoparametric hypersurfaces in real space forms with constant sectional curvature $c$ systematically [1]. Then many mathematicians started working on isoparametric functions and the classification of isoparametric hypersurfaces in real space forms [2, 3, 4]. If $f$ only satisfies the first equation of (1.1), $f$ is called a transnormal function.
@@
However, no proof was given in [15]. Later,  [14] supplied the proof of Theorem A and constructed an example to point out there is an essential difference between transnormal functions and isoparametric functions on hyperbolic spaces.
@@
$\left\{\begin{aligned} &F^{2}(\nabla f)=a(f),\\ &\Delta f=b(f).\end{aligned}\right$. (1.2)
@@
The fundamental structural result in [15] was generalized to Finsler geometry in [5]. Based on the structural result given in [5], in this paper, we try to generalize Theorem A to Finsler geometry. Unlike the Riemannian case, the geodesics are not necessarily reversible and the regularity of $f$, $a(t)$ and $b(t)$ in (1.2) are very poor, which causes more difficulties for the research.
For a connected Finsler manifold $(N,F)$, set $\overline{J}=f(N)$. Denote the focal variety of $f$ by $M_{\pm}$, where $M_{+}$ (resp. $M_{-}$) is the set that $f$ attains its global maximum (resp. minimum) value. Let $f$ be a $C^{2}$ transnormal function satisfying $F^{2}(\nabla f)=a(f)$, where $a(t)\neq 0$ for $t\in J$ and $a^{\prime}(t)\neq 0$ for $t\in\partial\overline{J}$. We called such $f$ an appropriate transnormal function. Let $g$ be the number of distinct principal curvatures for an isoparametric hypersurface. Then we have
@@
If $f$ is an appropriate transnormal function and f-f is a transnormal function on a Finsler space form $(N(c),F)$, then $f$ is isoparametric when $c\geq 0$, or $c<0$ and all principal curvatures $\kappa_{i}$ of a regular level hypersurface satisfying $|\kappa_{i}|\geq\sqrt{-c}$, $1\leq i\leq g$.
@@
Let $M_{\pm}$ be submanifolds and $f$ has no critical value in $J$. Then the condition in Theorem 1.1 that $f$ is an appropriate transnormal function can be weaken to $f$ is a transnormal function.
@@
Let $(N(c),F)$ be a forward Finsler space form and $r(x)=d_{F}(p,x)$ be a distance function from any $p\in N$. Then $f=\frac{1}{2}r^{2}(x)$ is a locally isoparametric function with $g=1$ on the cut-domain of $p$. Particularly, $f=-\cos(\sqrt{c}r)$ is a global isoparametric function on a standard Finsler sphere.
@@ 2.1 Finsler manifolds
$\nabla^{w}_{v}X(x)=\{v^{j}\frac{\partial X^{i}}{\partial x^{j}}(x)+\Gamma^{i}_{jk}(w)v^{j}X^{k}(x)\}\frac{\partial}{\partial x^{i}}.$ (2.3)
@@
where $X,Y\in T_{x}N$, $\nabla^{2}f(X)=\nabla^{\nabla f}_{X}\nabla f\in T_{x}N$ for $x\in N_{f}$. The Laplacian of $f$ can defined by
@@ 2.2 Anisotropic submanifolds
Let $(N,F)$ be an $n$-dimensional Finsler manifold and $\phi:M\to(N,F)$ be an (n1)(n-1)-dimensional immersion. Set $(x,y)\in TN$ with local coordinate $(x^{i},y^{i})$ and $(u,v)\in TM$ with local coordinate $(u^{a},v^{a})$, we have
@@
We call the eigenvalues of ${A}_{\textbf{n}}$, $k_{1},k_{2},\cdots,k_{m}$, the principal curvatures of $M$ with respect to n and $\hat{H}_{\textbf{n}}=\sum\limits_{a=1}^{m}k_{a}$ the anisotropic mean curvature of $M$.
[…]
@@ 2.3 Isoparametric functions and isoparametric hypersurfaces
More precisely, the function $f$ is called isoparametric (resp. $d\mu$-isoparametric) on $(N,F,d\mu)$, if (1.2) holds for $\Delta f=\hat{\Delta}f$ (resp. $\Delta f=\Delta_{\sigma}f$), which is defined by (2.6) (resp. (2.7)). If $(N,F,d\mu)$ has constant $\mathbf{S}$-curvature, then $f$ is isoparametric if and only if it is $d\mu$-isoparametric. For simplicity, we only consider $\Delta f$ as $\hat{\Delta}f$ in this paper. Obviously, all conclusions in this paper still hold for $d\mu$-isoparametric hypersurfaces if $\mathbf{S}$-curvature is a constant. Similar to Riemannian isoparametric hypersurfaces, we can give a geometric definition of isoparametric functions by anisotropic mean curvatures and principal curvatures.
Definition 2.1.
[9] A transnormal function $f$ is isoparametric on $(N,F)$ if and only if each regular level hypersurface of $f$ has constant anisotropic mean curvatures.
Definition 2.2.
[9] Let $(N,F)$ be a Finsler manifold with constant flag curvature. Then a transnormal function $f$ is isoparametric if and only if each regular level hypersurface of $f$ has constant principal curvatures.
Unlike Riemannian case, there exists some locally isoparametric hypersurfaces that cannot be extended to global isoparametric hypersurfaces (See Remark 4.3). Indeed, if some conditions are added, such extension can be achieved (See Remark 5.1). Hence, we need to give the definition of locally isoparametric functions on $(N,F)$. If there exist a neighborhood $V$ of $M$ in $(N,F)$ and an isoparametric function $f$ defined on $V$ such that $M$ is a regular level hypersurface of $f$, then $f$ is called a locally isoparametric function and $M$ is called a locally isoparametric hypersurface[13]. We can also give the geometric definition of locally isoparametric hypersurfaces in $(N,F)$.
Definition 2.3.
[7] A hypersurface $M$ in a Finsler manifold is a locally isoparametric hypersurface if and only if it and its sufficiently close parallel hypersurfaces have constant anisotropic mean curvatures.
Definition 2.4.
[13] $M$ is a locally isoparametric hypersurface in a Finsler manifold with constant flag curvature if and only if its principal curvatures are all constants.
3 Transnormal functions on Finsler space forms
@@ 3.1 Transnormal functions and focal varieties
For $\tau\in\mathcal{N}_{x}M$, define the normal exponential map $E:\mathcal{N}M\rightarrow N$ satisfying $E(x,\tau)=\exp_{x}\tau$, which is $C^{\infty}$ on $TN\setminus\{0\}$. The focal points of $M$ are the critical values of the normal exponential map $E$. The focal point $p\in N$ has multiplicity $m$ means that $E_{*}$ at $(x,\tau)$ has nullity $m>0$. The set of focal points is called a focal set, which is denoted by $\textmd{Foc}(f)$. The focal varieties of $f$ is denoted by $M_{\pm}$, which is the set where $f$ attains its global maximum value or global minimum value (if either of them exists). Denote $V=\{p\in N\big{|}\nabla f(p)=0\}$, which is the set of critical point of $f$. Indeed, $\textmd{Foc}(f)\subset M_{\pm}\subset V$.
Lemma 3.1.
[5] Let $f$ be a transnormal function on $(N,F)$. Set $\alpha,\beta]\subset\overline{J}$. If $x\notin f(V)$, where $x\in[\alpha,\beta)$, and $\beta\in f(V)$, then
$d(M_{\alpha},M_{\beta})=\int^{\beta}_{\alpha}\frac{df}{\sqrt{a(f)}}$ (3.1)
and the improper integral in (3.1) converges.
Remark 3.2.
In fact, from the proof of Lemma 3.1 in [5], we know that if $x\notin f(V)$, where $x\in(\alpha,\beta)$, and $\alpha,~\beta\in f(V)$, then (3.1) still holds.
Lemma 3.3.
[5] Let $(N,F)$ be a complete, connected Finsler manifold and $f$ be an appropriate transnormal function on $N$, then we have
(1)each focal variety of $f$ is a smooth submanifold in $N$.
@@
Let $f$ be an appropriate transnormal function on $(N,F)$. If $\gamma(s)$ is a maximal $f$-segment satisfying $\gamma(0)\in M_{+}$ (resp. $M_{-}$), then $\gamma^{\prime}(0)\in\mathcal{N}^{0}_{\gamma(0)}M_{+}$ (resp. $\mathcal{N}^{0}_{\gamma(0)}M_{-}$). Conversely, for any $\eta\in\mathcal{N}^{0}_{p}M_{+}$ (resp. $\mathcal{N}^{0}_{q}M_{-}$), there exists a maximal $f$-segment $\gamma(s)$ satisfying $\gamma(0)=p\in M_{+}$ (resp. $\gamma(0)=q\in M_{-}$) and $\gamma^{\prime}(0)=\eta$.
Remark 3.5.
If $f$ is only a transnormal function and $M_{\pm}$ are submanifolds, Lemma 3.4 holds.
For an appropriate transnormal function $f$ satisfying $F^{2}(\nabla f)=a(f)$, from Lemma 3.1, we define $d_{+}=d(M_{-},M_{+})=\int^{\beta}_{\alpha}\frac{dt}{\sqrt{a(t)}}$. Let f-f be a transnormal function satisfying $F^{2}(\nabla(-f))=\tilde{a}(-f)$. Due to $\nabla f=0$ if and only if $\nabla(-f)=0$, we know that f-f has no critical values in $f(N_{f})$. Since $M_{\pm}^{f}=M_{\mp}^{-f}$, we define $d_{-}=d(M_{+},M_{-})=\int^{\beta}_{\alpha}\frac{dt}{\sqrt{\tilde{a}(t)}}$. From Lemma 3.4, we immediately have
@@
Let $f$ be an appropriate transnormal function and f-f be a transnormal function on $(N,F)$. If $d_{+}<+\infty$ and $\gamma(s)$, $s\in[0,+\infty)$, is a unit speed geodesic such that $\gamma(s)\big{|}_{[0,d_{+}]}$ is a maximal $f$-segment, then $\gamma(s)\big{|}_{[d_{+},d_{+}+d_{-}]}$ is a maximal (f)(-f)-segment with respect to $F$.
@@
If $f$ is a transnormal function with no critical values in $J$ and $M_{\pm}$ are submanifolds, then Lemma 3.6 holds.
@@
In this case, $d_{+},d_{-}<+\infty$. From Lemma 3.8, for any $\kappa_{i}$, there exists some $s_{i}$ such that $\gamma(s_{i})\in Foc(f)$. Then From Lemma 3.6 and (3.5), there must exist $n_{i},m_{i}\in Z$ such that the principal curvatures can be expressed by
@@
From Lemma 3.8, for any $\kappa_{i}\neq 0$, there exists some $s_{i}$ such that $\gamma(s_{i})\in Foc(f)$.
@@
(2) $Foc(f)\neq\varnothing$. If $Foc(f)=M_{+}$, then $\kappa_{1}=\frac{1}{r}$. If $Foc(f)=M_{-}$ and $Foc(f)\neq M_{+}$, then $\kappa_{1}=\frac{1}{r+d_{-}}$. From [5], $p=\gamma(r_{0})$ is the focal point defined by the normal exponential map $\Phi(s,t):=E(\sigma(t),s\textbf{n}(\sigma(t)))$, where $\sigma(t)$ is a curve on $M$ and $\textbf{n}(\sigma(t))\in\mathcal{N}^{0}_{\sigma(t)}M$, if and only if $J(r_{0})=\frac{\partial\Phi}{\partial t}\big{|}_{t=0,s=r_{0}}=0$. From Cartan-Hadamard Theorem, $\gamma(s)$ passes through $Foc(f)$ once. Hence, from Lemma 3.6 and (3.5), we have $g=2$ and $\kappa_{2}=0$.
@@
From Cartan-Hadamard Theorem, there exists one $s_{i}$ at most such that $|\kappa_{i}|>\sqrt{-c}$. Similar to the analysis in Case 2, from Lemma 3.6 and (3.5), if $|\kappa_{i}|\geq\sqrt{-c}$, then the principal curvatures may be $\sqrt{-c}\coth\sqrt{-c}r$, $\sqrt{-c}\coth\sqrt{-c}(r+d_{-})$, $\sqrt{-c}$ or $\sqrt{-c}$, and $g\leq 2$.
Whether it is case 1, case 2 and case 3, the principal curvatures $\kappa_{i}$ of $M$ are all constants. From Definition 2.4, $M=f^{-1}(t_{0})$ is a locally isoparametric hypersurface and $f$ is a locally isoparametric function. Further more, from (2.11),
@@
which is an ODE with respect to an arc-length parameter $s$ and its initial values are determined by the constant principal curvatures of $M$. Combine Lemma 3.6 and (3.6), we know each level hypersurface of $f$ has constant principal curvatures. From Definition 2.2, $M=f^{-1}(t_{0})$ is a global isoparametric hypersurface and $f$ is a global isoparametric function. This completes the proof of Theorem 1.1. From the proof of Theorem 1.1, we immediately have
@@
Let $(N(c),F)$ be a Finsler space form. Suppose that $f$ is an appropriate transnormal function and f-f is a transnormal function on $N(c)$. If $c=0$, or $c<0$ and all principal curvatures $\kappa_{i}$ of a regular level hypersurface satisfying $|\kappa_{i}|\geq\sqrt{-c}$, $1\leq i\leq g$, then each regular level hypersurface has two distinct principal curvatures at most.
@@
In [12], we obtain the same result in different conditions. In fact, we need to rectify an omission of Theorems in [12]. That is, for $c<0$ and $c=0$, it also need to satisfy $|\kappa_{i}|>\sqrt{-c}$ and $\kappa_{i}\neq 0$, respectively.
@@
In local case, Theorem 3.9 doesn’t hold. That is, there exist locally isoparametric hypersurfaces in some Minkowski spaces with more than two principal curvatures (See Theorem 4.2 for detail).
If $F$ is reversible, $\tilde{a}(-f)=F^{*}(-df)=F^{*}(df)=F(\nabla f)=a(f)$. That is, $f$ is an appropriate transnormal function if and only if f-f is an appropriate transnormal function. Hence, we have
@@
4 Locally isoparametric hypersurfaces